# ROPA: Synthetic Robot Pose Generation for RGB-D Bimanual Data Augmentation

### v1

@@ I Introduction
footnotetext: *GSS holds concurrent appointments as a Professor at USC and as an Amazon Scholar. This paper describes work performed at USC and is not associated with Amazon.
@@ V-A Simulation Experiment Setup
We conduct simulation experiments using the open-source code from PerAct2 [18], which builds upon RLBench [60] to provide a benchmark for bimanual manipulation. We simplify certain tasks by reducing axes of variation (e.g., constraining the workspace) to enhance the performance of the Action Chunking with Transformers (ACT) baseline [19]. We use the following simulation tasks and success criteria:
  • Coordinated Lift Ball(CLB): lift a randomly spawned ball above $0.95\text{\,}\mathrm{m}$ in a $0.65\times 0.91$  $\text{\,}\mathrm{m}$ workspace (same as PerAct2).
  • Coordinated Lift Tray(CLT): lift a tray with an item above $1.2\text{\,}\mathrm{m}$ in a $0.46\times 0.64$  $\text{\,}\mathrm{m}$ workspace (70% of PerAct2’s original, no tray rotation).
  • Coordinated Push Box(CPB): push a large box (randomly rotated up to 4°) to a target area in a $0.59\times 0.82$  $\text{\,}\mathrm{m}$ workspace (90% of PerAct2’s original).
  • Bimanual Straighten Rope Easy(BSR): place both rope endpoints in their target areas in a $0.33\times 0.46$  $\text{\,}\mathrm{m}$ workspace (50% of PerAct’s original).
  • Coordinated Put Item in Drawer(CPID): place an item inside one of three specified drawers in a $0.46\times 0.64$  $\text{\,}\mathrm{m}$ workspace (70% of PerAct2’s original).
@@
We ensure fair comparisons by using identical training, validation, and test data with consistent environment seeds across all methods. Demonstrations are generated via RLBench’s waypoint-based motion planner. Each ACT policy trains on 100 episodes plus augmented data for 260,000 iterations, with checkpoints saved every 2,000 iterations. All checkpoints are validated on 25 episodes, and the best-performing checkpoint is evaluated on 25 unseen test episodes. We use the same ACT parameters as PerAct2. For real-world experiments, we use the final checkpoint with consistent starting configurations. To maintain consistency with our RGB-only evaluation policy, we adapt ACT to handle both RGB and depth modalities. We modify the input architecture by expanding from 3 to 6 channels.
For VISTA, we train ZeroNVS with VISTA’s default fine-tuning parameters.
@@ VI-A Real-World Experiment Setup
  • Lift Ball: lift a $0.35\text{\,}\mathrm{m}$ diameter ball above $0.25\text{\,}\mathrm{m}$ from randomized positions in a $0.64\times 0.20$  $\text{\,}\mathrm{m}$ region.
  • Push Block: push a $0.07\times 0.35\times 0.12$  $\text{\,}\mathrm{m}$ foam block (randomly placed and rotated up to 13°) past the front of a $0.97\times 0.43$  $\text{\,}\mathrm{m}$ workspace.
  • Lift Drawer: lift a $0.29\times 0.29\times 0.15$  $\text{\,}\mathrm{m}$ square drawer above $0.22\text{\,}\mathrm{m}$ from random positions/rotations (up to 25°) in a $0.48\times 0.38$  $\text{\,}\mathrm{m}$ region.
@@ VIII Conclusion
VIII-A Acknowledgments
Huge thanks to Charlene Yuen, Toan Nguyen, Haodi Hu, Ayano Hiranaka, Yunshuang Li, Jonathan Ong, and Vaidehi Dalmia for discussions and helpful writing feedback.

### v2

@@ I Introduction
footnotetext: All authors are with the Thomas Lord Department of Computer Science at the University of Southern California, USA. Correspondence: jchen567@usc.edu.
@@ V-A Simulation Experiment Setup
We conduct simulation experiments using the open-source code from PerAct2 [16], which builds upon RLBench [20] to provide a benchmark for bimanual manipulation. We simplify certain tasks by reducing axes of variation (e.g., constraining the workspace) to enhance the performance of the Action Chunking with Transformers (ACT) baseline [68]. We use the following simulation tasks:
  • Coordinated Lift Ball(CLB): Lift a randomly spawned ball.
  • Coordinated Lift Tray(CLT): Lift a tray with an item above.
  • Coordinated Push Box(CPB): Push a large box to a target area.
  • Bimanual Straighten Rope Easy(BSR): Place both rope endpoints in their target areas.
  • Coordinated Put Item in Drawer(CPID): Place an item inside one of three specified drawers.
@@
We ensure fair comparisons by using identical training, validation, and test data with consistent environment seeds across all methods. Demonstrations are generated via RLBench’s waypoint-based motion planner. To maintain consistency with our RGB-only evaluation policy, we adapt ACT to handle both RGB and depth modalities. We modify the input architecture by expanding from 3 to 6 channels. For VISTA, we train ZeroNVS with VISTA’s default fine-tuning parameters.
@@ VI-A Real-World Experiment Setup
  • Lift Ball: Lift a ball above a region.
  • Push Block: Push a block past the front of a workspace.
  • Lift Drawer: Lift a drawer above a region.
@@ VIII Conclusion
IX Acknowledgments
Huge thanks to SLURM Lab for discussions and helpful writing feedback.